OpenStudy (anonymous):

Solve the problem. Find all values of k so that the given points are \sqrt {29} units apart. (-5, 5), (k, 0) a. 3, 7 b. 7 c. -3, -7 d. -7

hero (hero):

just substitute values of k until you get it

hero (hero):

oh wait, it said all values of k, nevermind

OpenStudy (anonymous):

solve \[(k+5)^2+5^2=29\]

OpenStudy (anonymous):

\[(k+5)^2=4\] \[k+5=2\] or \[k+5=-2\] \[k=-3\] or \[k=-7\]

OpenStudy (anonymous):

The answwer is D

OpenStudy (anonymous):

I think too.

hero (hero):

Satellite, can you post a graphic display of this?

hero (hero):

I know it's distance formula but still...